### v2

# {textblock*}

Abstract
The remarkable performance achieved by Large Language Models (LLM) has driven research efforts to leverage them for a wide range of tasks and input modalities. In speech-to-text (S2T) tasks, the emerging solution consists of projecting the output of the encoder of a Speech Foundational Model (SFM) into the LLM embedding space through an adapter module. However, no work has yet investigated how much the downstream-task performance depends on each component (SFM, adapter, LLM) nor whether the best design of the adapter depends on the chosen SFM and LLM. To fill this gap, we evaluate the combination of 5 adapter modules, 2 LLMs (Mistral and Llama), and 2 SFMs (Whisper and SeamlessM4T) on two widespread S2T tasks, namely Automatic Speech Recognition and Speech Translation. Our results demonstrate that the SFM plays a pivotal role in downstream performance, while the adapter choice has moderate impact and depends on the SFM and LLM.
Index Terms:
automatic speech recognition, speech translation, LLM, foundation models, adapters
I Introduction
The success of Large Language Models (LLMs) [1] has attracted significant interest in extending their capabilities to handle various input modalities such as vision [2] and speech [3]. In the speech scenario, several studies [3, 4, 5, 6, 7] have proposed the integration of a pretrained Speech Foundation Model (SFM) encoder with a pretrained LLM through an adapter module, realizing the SFM+LLM new architectural paradigm [8]. The adapter can be decomposed into two components, as shown in Fig. 1: a length adapter, which compresses the speech sequence along the time dimension, and a modality adapter, which maps the compressed input into an embedding space compatible with the LLM. The SFM+LLM solution exploits, on the one hand, the SFM ability to extract high-quality semantic representations of the speech input and, on the other, the fluency and vast linguistic knowledge of LLMs, achieving competitive scores for widespread tasks such as Automatic Speech Recognition (ASR) [9] and Speech Translation (ST) [7].
However, being a recent trend, research efforts have mostly been devoted to demonstrating the effectiveness of this paradigm over traditional methods, without delving into the single design choices. Specifically, many architectural solutions have been proposed for the length adapter, often employed to both reduce the LLM computational costs and the modality mismatch with the textual sequences. These methods span from fixed downsampling, obtained either with a stack of strided convolutions [9] or with window-level Q-Former [3], to modules with variable compression rates that reduce the input sequence based on its semantic content, such as Continuous Integrate-and-Fire (CIF) [10] and CTC compression [11]. Nonetheless, a comprehensive study on the length adapter choice is missing: while some comparisons are present in the literature [5, 7], these evaluations are narrow in scope and assume the optimal solution is independent of the chosen SFM or LLM.
In this work, we explore whether a one-size-fits-all design choice for the length adapter exists that maximizes ASR and ST performance or if this depends on the selected SFM and LLM. Through a systematic comparison of a wide range of adapters proposed in the literature and by analyzing their impact in combination with widely used SFMs (Whisper [12] and SeamlessM4T [13]) and LLMs (Llama [1] and Mistral [14]), our contributions can be summarized as follows:
  • We prove that performance highly varies when a different SFM is selected (on average, more than 2 COMET points for ST and 1 WER for ASR), whereas the choice of the LLM and length adapter has a less pronounced impact on the final performance.
  • We show that there is no one-size-fits-all solution for the length adapter as its choice highly depends on the selected SFM and LLM combination.
Our codebase will be released under the Apache 2.0 License upon paper acceptance.
II Methodology
Fig. 1 illustrates the design of SFM+LLM solutions, highlighting the three key components: the SFM encoder, the adapter, and the LLM. The input audio $a$ is first processed by the SFM encoder $\mathcal{S}_{E}$, whose output goes through the adapter $\mathcal{A}$ and the resulting embeddings are concatenated to the embeddings $\mathcal{P}$ of a textual prompt. [^1] In our experiments, we use as textual prompt “can you transcribe $L_{s}$?” for ASR, where $L_{s}$ is the language of the audio, and “can you translate from $L_{s}$ to $L_{t}$?” for ST, where $L_{t}$ is the desired output language. The operation $J$, which joins the audio and prompt embeddings, is the same as that used in LLaVa [2] for prompting. Defining $\mathcal{L}$ as the LLM, the transcription or translation $y$ is then obtained as follows:
$y=\mathcal{L}\left(J\left(\mathcal{P},\mathcal{A}\left(\mathcal{S}_{E}(a)\right)\right)\right)$ (1)
As the main rationale behind the SFM+LLM solution comes from the possibility of training a high-quality ASR or ST system without large training datasets – thus with limited computational costs and memory requirements – we keep $\mathcal{L}$ and $\mathcal{S}_{E}$ frozen, training only $\mathcal{A}$. This solution is coherent with previous work [15], which showed that the gains obtained by fine-tuning the whole SFM encoder and LLM do not justify the additional costs. Within this framework, we answer our research question on the relative importance of the three components for the downstream performance by varying each of them as illustrated in the following sections.
II-A SFM Encoder
To investigate the impact of the SFM encoder, we use two widely recognized SFMs for speech representation extraction: Whisper [12] and SeamlessM4T [13]. In particular, we use the large version of both of them, namely Whisper large-v3, [^2] https://huggingface.co/openai/whisper-large-v3 and SeamlessM4T v2-large. [^3] https://huggingface.co/facebook/seamless-m4t-v2-large While Whisper is the most popular SFM in recent works on SFM+LLM [3, 4, 16, 17, 18], representing a natural choice, the usage of SeamlessM4T has never been explored to the best of our knowledge. Nonetheless, we opted for it not only for its recognized quality, but also because its design is very different from Whisper. SeamlessM4T is built with a customized version of Conformer layers [19] instead of Transformer ones, and also the compression factor of the input sequence is very different. While both process audio sequences where each vector of the sequence represents 10ms of audio, Whisper emits one vector every 20ms (2$\times$ downsampling), while SeamlessM4T encoder returns one vector every 160ms (16$\times$ downsampling). In light of these peculiarities, experimenting with these two SFM encoders lets us understand whether their behavior impacts on the best adapter design (in particular, the length adapter).
Figure 1: Schema of the SFM+LLM architecture.
II-B Adapter
As we keep the SFM and LLM frozen, we design adapters with high representation capacity, allowing for an effective mapping of the embeddings to the LLM input space. To do so, we follow [15, 4] that use a stack of vanilla Transformer [20] encoder layers with bidirectional self-attention as modality adapters, and investigate different methods as length adapters. The adapters are trained using a cross-entropy loss on the output of the LLM having the transcripts for ASR and translations for ST as target, unless stated otherwise. Overall, we investigate the following 5 types of adapters.
Base. 4 Transformer encoder layers are used and no length adaptation is performed.
Conv-based. 2 convolutional layers with stride 2 are introduced after the second layer of the Base adapter. No auxiliary loss is used. The final compression factor is 4.
CIF-based. Similarly to Conv-based, the adapter is extended by introducing a Continuous Integrate-and-Fire (CIF) [10] length adapter after the second Transformer layer. CIF is a sequence compression mechanism that accumulates input features over time and emits an output when a given integration threshold is reached, enabling variable-length sequence compression while preserving key information. To train this module, we add two auxiliary losses: a Connectionist Temporal Classification (CTC) loss [21] with the transcripts as target, following [22], and a quantity loss that controls the compression factor. The weight associated to both auxiliary losses is 0.1. On average, this corresponds to a compression factor of 2 with SeamlessM4T and 12 with Whisper.
CTC-based. In this case, the length adapter is a CTC-based compression [11], which collapses consecutive equal predictions of a CTC module by averaging the corresponding vectors, trained on the transcripts with an auxiliary CTC loss as done in CIF-based. On average, this corresponds to a compression factor of 1.5 for SeamlessM4T and 9 for Whisper.
WLQ-former. This adapter performs both modality and length adaptation with a window-level Q-Former [3]. This module processes variable-length encoded speech sequences by dividing them into fixed-length windows of encoded frames and feeding each of these non-overlapping windows to a Q-former architecture [23]. The Q-former uses a fixed and configurable number of learnable query vectors to attend to each window through cross-attention. As a result, the compression factor is controlled by the window length and the number of queries, which we set to 0.33 seconds and 1 respectively as per [3], and therefore results in 2 for SeamlessM4T and 16 for Whisper.
TABLE I: Compression rate for each configuration of SFM/Adapter.
SFM Adapter Compression ratio Sampling rate (Hz)
Seamless Base 1:1 6.25
CIF-based 3:1 2.08
Conv-based 4:1 1.56
CTC-based 2:1 3.12
WLQ-former 2:1 3.12
Whisper Base 1:1 50.00
CIF-based 25:1 2.00
Conv-based 4:1 12.50
CTC-based 13:1 3.85
WLQ-former 16:1 3.12
In Table I, we report the various compression factors for each adapter. We also distinguish between the SFMs (SeamlessM4T and Whisper) as their output sequence length is very different and this influences the compression factor, especially for the content-based adapters (CIF-based and CTC-based). For the content-based length adapters, we present the estimated compression ratios on the test set, evaluated after training. Notably, the CIF-based adapter reaches the highest compression ratio (25:1) with Whisper, which corresponds to a $\sim$2 Hz frequency of output vectors. The CTC-based, on the other hand, reduces the output sequence to $\sim$3 Hz using SeamlessM4T while it reduces the sequence to approximately 4 Hz with Whisper. The corresponding textual representations by the LLM tokenizers contain $\sim$2.7 token per second, i.e. slightly longer than those produced by CIF-based and slightly shorter than those produced by the CTC-based and WLQ-former.
II-C LLM Decoder
As LLMs, we select Mistral-7B-Instruct-v0.3, [^4] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 and Llama-3.1-8B-Instruct [^5] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct. The former is an English-centric model while the latter is trained in a multilingual setting, covering English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai. Moreover, Llama is the most popular choice in previous works [4, 24, 17]. Also in this case, our goal is to maximize the difference between the investigated LLMs.
III Experimental Settings
We trained all our models on CoVoST 2 [25] using English, German, Spanish, French, and Italian as source languages and German and English as target languages, and on MuST-C [26] using German, French, Italian, Spanish as target languages. The datasets for ASR were obtained by taking the audios in the source languages and their corresponding transcriptions for CoVoST, while for MuST-C we took audios and associated transcriptions from the English-German partition. We trained for 2 epochs, corresponding to 28k steps, on 4 NVIDIA GH200 96GB GPUs, using a micro-batch size of 10 samples and 4 gradient accumulation steps, for a total batch size of 160 samples, and took the last checkpoint as the final model. A single training run took approximately 24 hours. We used AdamW as the optimizer and a cosine scheduler for the learning rate with $10^{-4}$ as the peak value and a linear warmup of 840 steps. The hyperparameters of all the Transformer layers in the adapters are the same as in the BERTBASE model. [^6] https://huggingface.co/google-bert/bert-base-uncased The number of trainable parameters for each configuration is reported in Table II.
Results are reported as COMET [^7] Computed using https://huggingface.co/Unbabel/wmt22-comet-da. [27] and WER [^8] Computed using jiwer. for, respectively, ST and ASR on the CoVoST 2 test sets. Statistical significance is computed using bootstrap resampling [28].
TABLE II: Number of parameters for each adapter.
Adapter # trainable parameters (M)
len. adapt. mod. adapt. Total
Base 0 28.35 28.35
Conv-based 25.20 28.35 53.55
CIF-based 28.15 28.35 56.50
CTC-based 25.20 28.35 53.55
WLQ-former 33.09
TABLE III: ASR and ST results on CoVoST test sets. The best result for each (SFM, LLM) configuration is underlined, while the overall best is bolded. The difference with Base is statistically significant ($p<0.05$) unless for scores marked with .
SFM LLM Adapter ST - COMET ($\uparrow$) ASR - WER ($\downarrow$)
en-de de-en es-en fr-en it-en avg en es fr it de avg
SeamlessM4T Mistral Base 84.94 84.75 86.65 84.71 85.42 85.29 6.48 6.56 9.69 7.8 8.36 7.78
CIF-based 84.31 84.33 86.31 84.32 85.07 84.87 7.10 6.92 10.23 8.60 9.38 8.45
Conv-based 84.33 84.15 86.20 84.11 84.98 84.75 7.53 7.83 11.38 10.07 11.44 9.65
CTC-based 82.95 82.48 85.20 82.85 83.57 83.41 7.94 7.90 12.51 10.31 12.29 10.19
WLQ-former 84.67    84.71    86.60 84.59 85.29 85.17    6.38 6.80    9.83 8.05    8.48 7.91
Llama 3.1 Base 85.12 84.15 86.17 84.08 84.78 84.86 7.15 7.46 10.67 9.20 9.96 8.89
CIF-based 84.65 83.87 85.98 83.86 84.65 84.60 7.66    7.47 12.36 10.18 10.50 9.63
Conv-based 85.42 84.42 86.43 84.31 85.17 85.15    7.16 7.08 10.79 8.99    9.83 8.77
CTC-based 83.78 82.49 85.21 82.83 83.60 83.58 7.95 8.04 12.17 9.94 11.22 9.90
WLQ-former 85.65 84.84 86.66 84.68 85.39 85.44 6.62 6.69 9.96 7.97 8.71 7.99
Whisper Mistral Base 78.98 81.38 84.79 81.63 82.69 81.89 11.37 7.57 12.81 10.14 10.88 10.55
CIF-based 77.79 80.35 84.11 80.79 81.83 80.99 12.57 8.45 14.24 12.32 13.09 12.13
Conv-based 78.73 81.26    84.72    81.52 82.58 81.76 11.78 7.60 13.23 10.67 11.52 10.96
CTC-based 75.56 76.53 81.75 78.33 78.55 78.14 14.69 10.63 17.15 15.09 16.50 14.81
WLQ-former    79.07    81.44 84.92    81.68 82.92 82.00 11.82 8.21 13.60 15.77 12.55 12.39
Llama 3.1 Base 80.43 82.15 85.21 82.33 83.06 82.64 9.90 6.33 11.27 8.52 9.09 9.02
CIF-based 78.32 78.94 82.51 80.09 80.27 80.02 12.82 8.53 14.31 12.80 13.53 12.40
Conv-based 80.84 82.57 85.49 82.60 83.51 83.00    9.90 6.46    11.49    8.75    9.00 9.12
CTC-based 76.47 73.80 80.16 77.19 76.59 76.84 14.02 10.98 17.55 16.29 17.21 15.21
WLQ-former 79.95 81.56 84.88 81.56 82.89 82.17 11.98 7.90 14.52 11.10 12.84 11.67
IV Results
In Table III, we report the ASR and ST results of the SFM+LLM architectures with the various combinations of SFMs, LLMs, and length adapters introduced in Section II. As previous works mostly rely on BLEU [29], we highlight that our Whisper, Base adapter, and Llama model – a basic configuration adopted by previous SFM+LLM works – scores 28.7 BLEU on CoVoST en-de (the most popular language direction), a significantly higher result than the 25.1 BLEU of Qwen-Audio [30], one of the best performing SFM+LLM solutions. This confirms the soundness of our experimental settings and the competitiveness of our results.
First, we observe that the choice of the SFM is the most critical factor in terms of downstream performance. This is shown not only by the fact that, in each configuration, the version equipped with SeamlessM4T outperforms the counterpart with Whisper on both tasks (ASR and ST) on average, but also by the average improvement of more than 2 COMET on ST and more than 1 WER on ASR of the best SeamlessM4T configurations (with Llama and the WLQ-former or Base adapters) over the best Whisper ones (with Llama and the Conv-based adapter for ST and Llama and Base for ASR).
Instead, the choice of the LLM is less critical, as demonstrated by the small gap ($<$0.2 on both ASR and ST) between the best configuration with Llama (Seamless as SFM and WLQ-former as adapter) and that with Mistral (SeamlessM4T as SFM and Base as adapter). Moreover, the best results on average in ST are obtained with Llama, while the best ones in ASR with Mistral.
Second, results clearly show that there is no one-size-fits-all solution for the length adapter. Interestingly, the LLM plays an important role in the choice of the adapter. With Mistral, the Base adapter generally yields the best results, even though the WLQ-former is competitive, especially in ST. With Llama, instead, the best adapter varies with the SFM used. While WLQ-former is the best option with SeamlessM4T, the Conv-based and Base adapters emerge with Whisper, with the former being the best in ST and close in ASR, where differences among the two are almost always not statistically significant. Across all SFM and LLM configurations and tasks, the Base adapter always ranks first or second except for Llama+SeamlessM4T in ST, where it is third. Moreover, content-based length adapters consistently underperform other strategies. Together with the observation that there is no clear trend of the results with respect to the compression factor, these insights suggest that reducing the length mismatch between textual and speech representations is not critical for the quality of the outputs. However, reducing the speech sequence length lowers computational costs, making length adapters still a useful module to consider.
All in all, our results demonstrate the need for experimenting in different settings in terms of SFM and LLM when comparing adapter solutions, as improvements in one specific scenario may not generalize. In addition, LLMs show to be robust to input sequences of very different lengths, as the Base adapter, which does not compress the speech sequence, and the WLQ-former, which has high compression factors (16 with Whisper), achieve competitive scores in most settings.
V Conclusions
This work systematically analyzed the importance and design of the various building blocks that compose speech-to-text models by connecting an SFM encoder and an LLM through an adapter. To this aim, we compared all the combinations of 2 SFMs, 2 LLMs, and 5 adapters, which mostly differ for their length reduction module. With comprehensive experiments covering two tasks – ASR and ST – and 5 language directions, our results demonstrate that the choice of the SFM is the most critical factor influencing downstream performance. We also established that there is no one-size-fits-all solution for the length adapter, as the optimal choice varies depending on the specific combination of SFM and LLM. Notably, the Base and WLQ-former adapters, which feature very different compression factors, demonstrate strong performance across tasks, suggesting that reducing sequence length mismatch between speech and text is less crucial than previously assumed.
Acknowledgments
This paper has received funding from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People). We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid ACC Cyfronet AGH for providing computer facilities and support within computational grant no PLG/2024/016971. Marco Gaido was supported by the PNRR project FAIR - Future AI Research (PE00000013), under the NRRP MUR program funded by the NextGenerationEU. This work has been carried out while Pierfrancesco Melucci was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with Università di Roma Tre.